ALX Corp. has the following financial statements for years 2018 and 2017:

$m $m
Income statement 2018 2017
Revenue $          1,500 $          1,300
EBITDA $              160 $              150
Dep. & Amor $                30 $                25
EBIT $              130 $              125
Net Income $                21 $                20
EBT $              109 $              105
Taxes (35%)
Net Income
Balance Sheet 2018 2017
Current Assets $              340 $              250
Net Plant & Equip. $              370 $              340
Total Assets $              710 $              690
Current Liabilities $              260 $              240
LT Debt $              100 $              100
Equity & Ret. Earnings $              350 $              350
Total Liab. & Equity $              710 $              690

The 2018 Free Cash Flows (FCFs) for 2018 are:

Select one:

a. $25m

b. $145m

c. None of he Above

d. $115m

e. $55m
Answer #1

FCF = (EBITDA + Depreciation - Change in Net working Capital - capital expenditure)*65%

= (160 + 30 - ((CA-CL in 2018) - (CA-CL in 2018)) - (plant and equipment in 2018 - plant and equipment in 2017))*65%

= (190 -((340-260)-(250-240))-(370-340))**65% = 58.5

C) None of the above